OpenStudy (anonymous):

x^2 - 6x = 8 complete the square

OpenStudy (anonymous):

To complete the square, add 9 to both sides x^2 - 6x + 9 = 8 + 9 it can be written as (x)^2 - 2(x)(3) + (3)^2 = 17 LHS becomes the complete square of (x - 3)^2 so (x - 3)^2 = 17 taking root of both sides x - 3 = ± root(17) x = 3 ± root (17)